OpenStudy (anonymous):

the equation sin(pi/6-x) is equal to______?

OpenStudy (mathmale):

And so we meet again!! Have you considered applying the identity sin (a-b) = sin a cos b - cos a sin b?

OpenStudy (anonymous):

i dont quite know how to use that, walk me through it? maybe then we can reevaluate our trade.

OpenStudy (mathmale):

I'd thought our trade was cast in stone. :( Hints: sin pi/2 = 1; cos pi/2 = 0 Starting with \[\sin (a-b) = \sin a \cos b - \cos a \sin b\] write pi/2 in place of a and x in place of b. You up to doing that?

OpenStudy (anonymous):

noooooooo

OpenStudy (anonymous):

that seems so tedious